er 03 04 Total
Diane Buswell is preparing the 2020 budget for one of Current Designs' rotomolded kayaks. Extensive meetings with members of the sales department and executive team have Quarter 1 Quarter 2 Quarter 3 Quarter 4 1,900 kayaks 2,600 kayaks 1,550 kayaks 1,550 kayaks Current Designs' polRY to have finished goods ending inventory in a quarter equal to 15% of the next quarter's anticipated sales, preliminary sales project ons for 2021 are 1,100 units for the first quarter and 2,600 units for the second quarter. Ending inventory of finished goods at December 31, 2019, will be 285 rotomolded kayaks. powder should be 25% of the amount needed for production in the next quarter. Assume that the ending inventory of polyethylene powder on December 31, 2019, is 19,800 pounds. The finishing kits can be assembled as they are needed. As a result, Current Designs does not maintain a significant inventory of the finishing kits. powder used in these kayaks costs $1.20 per pound, and the finishing kits cost $160 each. Production of a single kayak requires 4 hours of time by more experienced, type I employees and 5 hours of finishing time by type II employees. The type I employees are paid $18 per hour, and the type II employees are paid $15 per hour. ng and administrative expenses for this lne are expected to be $42 per unit sold plus soooo per quarter. Manufacturing overhead is assigned at 160% of labor costs.
er 03 04 Total
Answer #1

CURRENT DESIGNS

Production Budget

For the Year Ending December 31, 2020

Quarter 1

Quarter 2

Quarter 3

Quarter 4

Total

Expected Unit Sales

1900

2600

1550

1550

7600

+ Desired Ending Finished Goods

390

233

233

165

165

Total Required Units

2290

2833

1783

1715

7765

- Beginning Finished Goods Units

285

390

233

233

285

Required Production Units

2005

2443

1550

1482

7480

2600*15% =390

1550*15% =233

1100*15% =

CURRENT DESIGNS

Direct materials Budget

For the Year Ending December 31, 2020

Quarter 1

Quarter 2

Quarter 3

Quarter 4

Total

Units to be produced

2005

2443

1550

1482

7480

Pounds of polyethylene powder per unit

44

44

44

44

44

Total pounds needed for production

88220

107492

68200

65208

329120

Add: Desired ending inventory

26873

17050

16302

14575

14575

Total pounds of powder required

115093

124542

84502

79783

343695

Less: Beginning inventory of powder

19800

26873

17050

16302

19800

Pounds of polyethylene powder to be purchased

95293

97669

67452

63481

323895

Cost per Pound

1.20

1.20

1.20

1.20

1.20

Cost of polyethylene powder to order

114351.60

117202.80

80942.40

76177.20

388674.00

Cost of required finishing kits

320800

390880

248000

237120

1196800

Total costs for direct materials

435151.60

508082.80

328942.40

313297.20

1585474

Total units to be produced in the first quarter of 2021 = 1100+(2600*15%)-165 = 1325

Total pounds needed for production = 1325*44 = 58300